# Light of Normals: Unified Feature Representation for Universal Photometric Stereo

### v2

@@ 5 Conclusion
Firstly, we design the light registration tokens to improve the handling of global illumination. While lighting information predominantly exhibits global characteristics across multi-light inputs [25, 28, 31], traditional attention mechanisms in existing Universal Photometric Stereo (PS) methods often fail to fully leverage this distributed information. This deficiency can hinder effective illumination-geometry separation, motivating our specialized token-based strategy. Drawing inspiration from advancements like [9], and further considering the inherent illumination-dependency of the Universal PS task, we introduce these light registration tokens to explicitly capture and represent decoupled global lighting information within our framework.
To facilitate the perception of distinct lighting components, we additionally introduce three specialized light tokens: $x_{\text{hdri}}\in\mathbb{R}^{1\times C}$, designated for perceiving spatially-varying environment light; $x_{\text{point}}\in\mathbb{R}^{1\times C}$, tailored for point lights (which often contribute high-frequency illumination effects); and $x_{\text{area}}\in\mathbb{R}^{1\times C}$, for directional light (typically representing low-frequency illumination sources). Subsequently, this set of three specialized light tokens is prepended to the token sequences derived from $F_{s,f}^{d}$ (features from the downsampled image stream) and $F_{s,f}^{w}$ (features from the wavelet components stream), respectively, leading to:$F_{s,f,r}^{d}\in\mathbb{R}^{L^{\prime}\times D}$ and $F_{s,f,r}^{w}\in\mathbb{R}^{4\times L^{\prime}\times D}$, where $L^{\prime}=L+3$.
@@
Then $F_{s,f,r}^{d}$ and $F_{s,f,r}^{w}$ are fed into our Interleaved Attention block [53] to enhance inter-intra feature communication. Specifically, our alternating attention block contains four attention layers, which can be represented as: frame $\rightarrow$ light-axis $\rightarrow$ global $\rightarrow$ light-axis.
@@
Previous work has found that feature communication within the encoder is very important [28, 31, 20], but their methods are often limited to patch-level local light-axis attention. Our alternating attention, however, breaks such limitations. On the one hand, it incorporates frame attention to enhance inter-image communication. On the other hand, we have added global attention, a more comprehensive global operation, allowing intra-image features to also be extended from the patch level to the image level. Upon processing by these four cascaded Interleaved Attention blocks (the number chosen to maintain a manageable parameter count, although more blocks could potentially be employed), we obtain the deep feature representations denoted as $F_{d,f,r}^{d}\in\mathbb{R}^{L^{\prime}\times D}$ and $F_{d,f,r}^{w}\in\mathbb{R}^{4\times L^{\prime}\times D}$. It is worth noting that the attention operations inherent in these blocks do not alter the fundamental shapes of these tensor sequences.
@@
To ensure the light registers tokens $x_{\text{hdri}}$, $x_{\text{point}}$, $x_{\text{area}}$ effectively capture global illumination, as direct supervision of the decoded light map is challenging, we introduce a light-aware feature alignment strategy during training  [60, 59, 49]. Since we train LINO-UniPS on our own rendered synthetic dataset PS-Verse, for every scene within this training dataset, we have access to its corresponding lighting information from the rendering process. This includes: the environment light HDRI map, the directions and intensities of point lights, and the direction and intensity of the directional light. Mathematically, these are denoted as $L_{\text{hdri}}\in\mathbb{R}^{H\times W\times 3}$, $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 4}$, and $L_{\text{area}}\in\mathbb{R}^{M_{2}\times 5}$, respectively. For $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 4}$, where $M_{1}$ is the number of point lights, its first three dimensions denote position and the fourth denotes intensity. For the directional light component $L_{\text{area}}\in\mathbb{R}^{M_{2}\times 5}$, where $M_{2}$ denotes the number of directional lights, its first three dimensions specify position, the fourth denotes the directional light size, and the fifth indicates intensity. Subsequently, we encode the lighting components $L_{\text{hdri}}$, $L_{\text{point}}$, and $L_{\text{area}}$ by projecting them into a $D$-dimensional feature space, thereby obtaining their respective representations $\mathbf{l}^{h}_{\text{hdri}}\in\mathbb{R}^{D}$, $\mathbf{l}^{h}_{\text{point}}\in\mathbb{R}^{D}$, and $\mathbf{l}^{h}_{\text{area}}\in\mathbb{R}^{D}$. Similarly, the light tokens $x_{\text{env}}$, $x_{\text{point}}$, and $x_{\text{dir}}$, after being processed by the cascaded alternating attention blocks, are projected into the same $D$-dimensional feature space. This yields their respective high-dimensional representations: $\mathbf{x}^{h}_{\text{hdri}}$, $\mathbf{x}^{h}_{\text{point}}$, and $\mathbf{x}^{h}_{\text{area}}$, all in $\mathbb{R}^{D}$. Specifically, this projection is realized using three structurally similar two-layer Multi-Layer Perceptrons (MLPs). Within this common embedding space, we employ cosine similarity to supervise and align their respective feature distributions. Cosine similarity is chosen as it effectively measures the directional concordance between feature vectors, making it suitable for aligning representations of different lighting characteristics. This supervision translates into three distinct loss functions, denoted as $\mathcal{L}_{\text{hdri}}$, $\mathcal{L}_{\text{point}}$, and $\mathcal{L}_{\text{area}}$. Their respective mathematical formulations are:
$\mathcal{L}_{\text{hdri}}$ $=1-\sum(\mathbf{l}_{\text{hdri}}^{h}\cdot\mathbf{x}_{\text{hdri}}^{h})$ (A1a)
@@
$\mathcal{L}_{\text{area}}$ $=1-\sum(\mathbf{l}_{\text{area}}^{h}\cdot\mathbf{x}_{\text{area}}^{h})$ (A1c)
A1.1.4 DPT-Based Fusion Module:
@@
The subsequent discussion details the fusion module applied to features extracted during the initial stages of our Encoder. The overall feature fusion process is DPT-based [45]. For clarity, we will first illustrate this process using the features derived from the downsampled image features $F_{d,f,r}^{d}$ as the primary example. Within each Interleaved Attention module of the encoder, features obtained from its four internal attention—frame, light-axis, global, and ight-axis—are first concatenated along the feature dimension. This operation yields an aggregated feature set for each module, denoted as $F_{d,f}^{d,(i)}\in\mathbb{R}^{L\times 4D}$, where $i\in\{1,2,3,4\}$ represents the index of the $i$-th attention block. It is crucial to note that the three additional light registration tokens (introduced previously) do not participate in this specific feature aggregation (concatenation) process. Then, the aggregated features $F_{d,f}^{d,(i)}$ from four selected Interleaved Attention blocks then undergo a series of projection and resizing operations. These operations transform them into a four-level feature pyramid, with hierarchical features $H_{1},H_{2},H_{3},$ and $H_{4}$. Their respective shapes are $F\times C\times H_{0}/2\times W_{0}/2$, $F\times 2C\times H_{0}/4\times W_{0}/4$, $F\times 4C\times H_{0}/8\times W_{0}/8$, and $F\times 4C\times H_{0}/16\times W_{0}/16$. In this context, $F$ represents the number of multi-light input images, $C$ is a base feature channel dimension (256), and $H_{0},W_{0}$ refer to the spatial resolution of the original, full-sized input images. These four levels of hierarchical features ($H_{1}$ through $H_{4}$) are subsequently progressively fused using residual convolutional blocks [22], ultimately producing a fused feature representation $F_{\text{fused}}^{d}\in\mathbb{R}^{F\times\frac{H}{2}\times\frac{W}{2}\times C}$. A similar fusion process is applied to the features derived from the wavelet components path ($F_{d,f,r}^{w}$ ), yielding a corresponding fused representation, $F_{\text{fused}}^{w}\in\mathbb{R}^{4\times F\times\frac{H}{2}\times\frac{W}{2}\times C}$. It is noteworthy that this strategy of selecting four feature levels for hierarchical fusion can be adapted if more than four Interleaved Attention blocks are employed in the encoder’s initial stages. For instance, if six such blocks are utilized, features from blocks indexed 1, 2, 4, and 6 might be selected to form the pyramid. Similarly, for an eight-block configuration, features from blocks 1, 3, 5, and 7 could be chosen as inputs to the hierarchical fusion pathway.
A1.1.5 WaveUpSample:
@@
To obtain the final encoder output $F_{\text{enc}}\in\mathbb{R}^{F\times H\times W\times C}$, the features derived from the downsampled image path $F_{\text{fused}}^{d}$ and those from the wavelet components path $F_{\text{fused}}^{w}$ should be integrated. The process is as follows: first, the feature map $F_{\text{fused}}^{d}$ is upsampled, yielding an representation $F_{\text{fused}}^{\text{up}}\in\mathbb{R}^{F\times H\times W\times C}$. Concurrently, for $F_{\text{fused}}^{w}$, which originates from the wavelet-transformed inputs, an inverse wavelet transform is applied to convert it back to the spatial domain, resulting in $F_{\text{fused}}^{\text{dwt}}\in\mathbb{R}^{F\times H\times W\times C}$. Finally, these two processed feature sets, $F_{\text{fused}}^{up}$ and $F_{\text{fused}}^{dwt}$, are element-wise summed. A Gaussian blur is subsequently applied to this sum to promote a smoother and more effective fusion of these potentially cross-domain features, ultimately producing the final encoder representation $F_{\text{enc}}\in\mathbb{R}^{F\times H\times W\times C}$.
@@
The decoder architecture in our LINO-UniPS is largely identical to that of SDM-UniPS [31]; for clarity, we briefly outline its key components and rationale here.
@@
A common initial step in PS for surface normal estimation is the pixel-wise aggregation of spatial-light features along the illumination axis, effectively reducing $F$ light channels to a single representation per pixel using input images $I_{f}$ and their corresponding encoded features $F_{\text{enc}}^{f}$. We introduce an approach, termed the pixel-sampling Transformer [58, 8, 57, 56], which uniquely operates on a fixed count ($m$, e.g., $m=2048$) of randomly chosen pixel locations. This strategy offers distinct advantages: it maintains a constant memory footprint per sample set regardless of image dimensions, thus ensuring excellent scalability; furthermore, by processing a sparse, randomly distributed set of points, it substantially curtails over-smoothing artifacts often prevalent in dense convolutional operations. The practical implementation of the pixel-sampling Transformer involves selecting $m$ random pixels, denoted $\{x_{i}\}_{i=1}^{m}$, from the valid (masked) region of the input image. For each sampled pixel $x_{i}$, its associated features $F_{\text{enc}}^{f}(x_{i})\in\mathbb{R}^{C}$ are obtained. These features $F_{\text{enc}}^{f}(x_{i})$ are then combined through element-wise addition with $I^{\prime}_{f}(x_{i})\in\mathbb{R}^{C}$. The term $I^{\prime}_{f}(x_{i})$ represents a high-dimensional projection of the raw pixel observations $I_{f}(x_{i})\in\mathbb{R}^{3}$, and this projection is performed by a two-layer MLP with the objective of enhancing the representational power of these raw observations by mapping them to this higher-dimensional space. Notably, our strategy of first projecting the raw observations to $\mathbb{R}^{C}$ and then performing addition differs from SDM-UniPS [31], which typically employ direct concatenation of features and the raw observations. These added per-pixel features $F_{\text{add}}^{f}(x_{i})$ are subsequently condensed into compact descriptors $A(x_{i})$ by employing Pooling by Multi-head Attention (PMA) [35]. The resulting collection of $m$ descriptors, $A(x_{i})_{i=1}^{m}$, is then fed into a Transformer network. The processing for each of the $m$ sampled points involves applying frame attention and light-axis attention to aggregate non-local context and cross-image information. Following this, a two-layer MLP is utilized to predict the surface normal vector for each location. These sparsely predicted normals are then systematically merged—for instance, through spatial interpolation or a dedicated upsampling module—to reconstruct the full-resolution surface normal map corresponding to the original input image dimensions. In essence, the pixel-sampling Transformer facilitates the modeling of robust non-local dependencies with notable computational efficiency, while concurrently preserving fine details in the output normal map. This makes the approach particularly well-suited for physics-based vision tasks that involve high-resolution imagery.
@@
In this part, we elaborate on the composition of our total training loss and the design of the respective weights for its constituent components. Based on the definitions provided in Eq. 4, Eq. 3.1 and Eq. A1, the overall training loss $\mathcal{L}$ for our LiNO-UniPS method can be expressed as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{hdri}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{area}}+\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$
Let us define the confidence-weighted normal reconstruction loss as $\mathcal{L}_{\text{conf}}$ and the normal gradient supervision loss as $\mathcal{L}_{\text{gradient}}$. The overall training loss $\mathcal{L}$ can then be expressed as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{hdri}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{area}}+\lambda_{4}\mathcal{L}_{\text{conf}}+\lambda_{5}\mathcal{L}_{\text{gradient}},$
@@
A primary motivation for our LiNO-UniPS framework is to advance beyond prior Universal PS methods by specifically improving the handling of challenging high-frequency regions. We identify these regions based on large magnitudes of the surface normal gradients, as these directly reflect geometric complexity. We deliberately avoid using gradients derived from the input RGB multi-light images as the primary criterion for this identification. The rationale is that while RGB gradients are indeed large in areas of intricate geometric detail, they can also exhibit high magnitudes in regions with significant basecolor variations, which do not necessarily correspond to the geometric high-frequency features we aim to emphasize and reconstruct accurately. Consequently, our methodology incorporates a loss function that is directly informed by surface normal gradients, rather than relying on a naive MSE.
@@
$\lambda_{1}=\frac{0.1}{(\mathcal{L}_{\text{hdri}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{2}=\frac{0.1}{(\mathcal{L}_{\text{point}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{3}=\frac{0.1}{(\mathcal{L}_{\text{area}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{4}=1,\lambda_{5}=\frac{0.1}{(\mathcal{L}_{\text{gradient}}/\mathcal{L}_{\text{conf}})_{\text{sg}}}$
@@
Our decision to set $\lambda_{4}$ to $1$ is because $\mathcal{L}_{\text{conf}}$ serves as the principal component in our overall loss function. The remaining auxiliary losses are then scaled using the adaptive weighting mechanism detailed in Eq. A1.1.7. This mechanism constrains their magnitudes to $0.1$ times that of the primary loss’s detached value, $(\mathcal{L}_{\text{conf}})_{\text{sg}}$, while still allowing their gradients to backpropagate fully. Such a strategy effectively positions these auxiliary losses to act as regularizers to the main learning objective, rather than allowing disparate loss magnitudes to vie for dominance and potentially destabilize training. This controlled weighting is crucial for ensuring stable and efficient training of LiNO-UniPS, mitigating issues such as excessively slow convergence or even training failure that can arise from an unbalanced multi-term loss function.
A1.2 Dataset Analysis and Presentation

### v4

@@ 5 Conclusion
Firstly, we design the Light Register Tokens to improve the handling of global illumination. While lighting information predominantly exhibits global characteristics across multi-light inputs Ikehata (2018; 2022; 2023), traditional attention mechanisms in existing Universal PS methods often fail to fully leverage this distributed information. This deficiency can hinder effective illumination-geometry separation, motivating our specialized token-based strategy. Drawing inspiration from advancements like Darcet et al. (2024), and further considering the inherent illumination-dependency of the Universal PS task, we introduce these Light Register Tokens to explicitly capture and represent decoupled global lighting information within our framework.
[…]
@@
Then $F_{s,f,r}^{d}$ and $F_{s,f,r}^{w}$ are fed into our Interleaved Attention Block Wang et al. (2025) to enhance inter-intra feature communication. Specifically, our Interleaved Attention Block contains four attention layers, which can be represented as: frame $\rightarrow$ light-axis $\rightarrow$ global $\rightarrow$ light-axis.
@@
Previous work has found that feature communication within the encoder is very important Ikehata (2022; 2023); Hardy et al. (2024), but their methods are often limited to patch-level local light-axis attention. Our alternating attention, however, breaks such limitations. On the one hand, it incorporates frame attention to enhance inter-image communication. On the other hand, we have added global attention, a more comprehensive global operation, allowing intra-image features to also be extended from the patch level to the image level. Upon processing by these four cascaded Interleaved Attention Blocks (the number chosen to maintain a manageable parameter count, although more blocks could potentially be employed), we obtain the deep feature representations denoted as $F_{d,f,r}^{d}\in\mathbb{R}^{L^{\prime}\times D}$ and $F_{d,f,r}^{w}\in\mathbb{R}^{4\times L^{\prime}\times D}$. It is worth noting that the attention operations inherent in these blocks do not alter the fundamental shapes of these tensor sequences.
@@
To ensure the Light Register Tokens $x_{\text{env}}$, $x_{\text{point}}$, $x_{\text{direction}}$ effectively capture global illumination, as direct supervision of the decoded light map is challenging, we introduce a light-aware feature alignment strategy during training Yu et al. (2025); Yao et al. (2025); Song et al. (2025). Since we train LINO UniPS on our own rendered synthetic dataset PS-Verse, for every scene within this training dataset, we have access to its corresponding lighting information from the rendering process. This includes: the HDRI environment map, the positions, the distance to camera and intensities of point lights, and the positions, the distance to camera, intensities and areas of direction lights. Mathematically, these are denoted as $L_{\text{env}}\in\mathbb{R}^{H\times W\times 3}$, $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 5}$, and $L_{\text{direction}}\in\mathbb{R}^{M_{2}\times 6}$, respectively. For $L_{\text{point}}\in\mathbb{R}^{M_{1}\times 5}$, where $M_{1}$ is the number of point lights, its first three dimensions denote position, the fourth denotes the distance and the fifth denotes the intensity. For the directional light component $L_{\text{direction}}\in\mathbb{R}^{M_{2}\times 6}$, where $M_{2}$ denotes the number of directional lights, its first three dimensions specify position, the fourth denotes the distance, the fifth denotes the directional light size, and the sixth indicates intensity. Subsequently, we encode the lighting components $L_{\text{env}}$, $L_{\text{point}}$, and $L_{\text{direction}}$ by projecting them into a $D$-dimensional feature space, thereby obtaining their respective representations $\mathbf{l}^{h}_{\text{env}}\in\mathbb{R}^{D}$, $\mathbf{l}^{h}_{\text{point}}\in\mathbb{R}^{D}$, and $\mathbf{l}^{h}_{\text{direction}}\in\mathbb{R}^{D}$. Similarly, the light tokens $x_{\text{env}}$, $x_{\text{point}}$, and $x_{\text{direction}}$, after being processed by the cascaded Interleaved Attention Blocks, are projected into the same $D$-dimensional feature space. This yields their respective high-dimensional representations: $\mathbf{x}^{h}_{\text{env}}$, $\mathbf{x}^{h}_{\text{point}}$, and $\mathbf{x}^{h}_{\text{direction}}$, all in $\mathbb{R}^{D}$. Specifically, this projection is realized using three structurally similar two-layer Multi-Layer Perceptrons (MLPs). Within this common embedding space, we employ cosine similarity to supervise and align their respective feature distributions. Cosine similarity is chosen as it effectively measures the directional concordance between feature vectors, making it suitable for aligning representations of different lighting characteristics. This supervision translates into three distinct loss functions, denoted as $\mathcal{L}_{\text{env}}$, $\mathcal{L}_{\text{point}}$, and $\mathcal{L}_{\text{direction}}$. Their respective mathematical formulations are:
$\mathcal{L}_{\text{env}}$ $=1-\sum(\mathbf{l}_{\text{env}}^{h}\cdot\mathbf{x}_{\text{env}}^{h})$ (A1a)
@@
$\mathcal{L}_{\text{direction}}$ $=1-\sum(\mathbf{l}_{\text{direction}}^{h}\cdot\mathbf{x}_{\text{direction}}^{h})$ (A1c)
Table A1: Comparison of PBR material prediction performance. A dash (-) indicates the method does not provide the corresponding output.
Method Normal Albedo Metallic Roughness
[…]
UniPS Ikehata (2022) 24.67 - - - - - -
IDarb Li et al. (2025) 29.51 25.11 0.9271 25.62 0.9148 24.85 0.9342
SDM UniPS Ikehata (2023) 10.25 24.04 0.9076 23.68 0.9060 23.87 0.9289
Uni MS-PS Hardy et al. (2024) 8.92 - - - - - -
Ours 4.51 30.49 0.9529 32.99 0.9357 31.32 0.9670
Figure A1: Qualitative comparison on PSVerse Testdata. LINO UniPS demonstrates superior PBR estimation compared to all other methods.
A1.1.4 Fusion Block:
@@
The subsequent discussion details the Fusion Block applied to features extracted during the initial stages of our Encoder. The overall feature fusion process is DPT-based Ranftl et al. (2021). For clarity, we will first illustrate this process using the features derived from the downsample image features $F_{d,f,r}^{d}$ as the primary example. Within each Interleaved Attention module of the encoder, features obtained from its four internal attention—frame, light-axis, global, and light-axis—are first concatenated along the feature dimension. This operation yields an aggregated feature set for each module, denoted as $F_{d,f}^{d,(i)}\in\mathbb{R}^{L\times 4D}$, where $i\in\{1,2,3,4\}$ represents the index of the $i$-th attention block. It is crucial to note that the three additional Light Register Tokens (introduced previously) do not participate in this specific feature aggregation (concatenation) process. Then, the aggregated features $F_{d,f}^{d,(i)}$ from four selected Interleaved Attention Blocks then undergo a series of projection and resizing operations. These operations transform them into a four-level feature pyramid, with hierarchical features $H_{1},H_{2},H_{3},$ and $H_{4}$. Their respective shapes are $F\times C\times H_{0}/2\times W_{0}/2$, $F\times 2C\times H_{0}/4\times W_{0}/4$, $F\times 4C\times H_{0}/8\times W_{0}/8$, and $F\times 4C\times H_{0}/16\times W_{0}/16$. In this context, $F$ represents the number of multi-light input images, $C$ is a base feature channel dimension (256), and $H_{0},W_{0}$ refer to the spatial resolution of the original, full-sized input images. These four levels of hierarchical features ($H_{1}$ through $H_{4}$) are subsequently progressively fused using residual convolutional blocks He et al. (2015), ultimately producing a fused feature representation $F_{\text{fused}}^{d}\in\mathbb{R}^{F\times\frac{H}{2}\times\frac{W}{2}\times C}$. A similar fusion process is applied to the features derived from the wavelet components path ($F_{d,f,r}^{w}$ ), yielding a corresponding fused representation, $F_{\text{fused}}^{w}\in\mathbb{R}^{4\times F\times\frac{H}{2}\times\frac{W}{2}\times C}$. It is noteworthy that this strategy of selecting four feature levels for hierarchical fusion can be adapted if more than four Interleaved Attention Blocks are employed in the encoder’s initial stages. For instance, if six such blocks are utilized, features from blocks indexed 1, 2, 4, and 6 might be selected to form the pyramid. Similarly, for an eight-block configuration, features from blocks 1, 3, 5, and 7 could be chosen as inputs to the hierarchical fusion pathway.
A1.1.5 Wavelet UpSample Module:
@@
To obtain the final encoder output $F_{\text{enc}}\in\mathbb{R}^{F\times H\times W\times C}$, the features derived from the downsample image path $F_{\text{fused}}^{d}$ and those from the wavelet components path $F_{\text{fused}}^{w}$ should be integrated. The process is as follows: first, the feature map $F_{\text{fused}}^{d}$ is upsampled, yielding a representation $F_{\text{fused}}^{\text{up}}\in\mathbb{R}^{F\times H\times W\times C}$. Concurrently, for $F_{\text{fused}}^{w}$, which originates from the wavelet-transformed inputs, an inverse wavelet transform is applied to convert it back to the spatial domain, resulting in $F_{\text{fused}}^{\text{dwt}}\in\mathbb{R}^{F\times H\times W\times C}$. Finally, these two processed feature sets, $F_{\text{fused}}^{up}$ and $F_{\text{fused}}^{dwt}$, are element-wise summed. A Gaussian blur is subsequently applied to this sum to promote a smoother and more effective fusion of these potentially cross-domain features, ultimately producing the final encoder representation $F_{\text{enc}}\in\mathbb{R}^{F\times H\times W\times C}$.
@@
The decoder architecture in our LINO UniPS is largely identical to that of SDM UniPS Ikehata (2023); for clarity, we briefly outline its key components and rationale here.
@@
A common initial step in PS for surface normal estimation is the pixel-wise aggregation of spatial-light features along the illumination axis, effectively reducing $F$ light channels to a single representation per pixel using input images $I_{f}$ and their corresponding encoded features $F_{\text{enc}}^{f}$. We introduce an approach, termed the pixel-sampling Transformer Wu et al. (2024b); Contributors (2023); Wu et al. (2022; 2024a), which uniquely operates on a fixed count ($m$, e.g., $m=2048$) of randomly chosen pixel locations. This strategy offers distinct advantages: it maintains a constant memory footprint per sample set regardless of image dimensions, thus ensuring excellent scalability; furthermore, by processing a sparse, randomly distributed set of points, it substantially curtails over-smoothing artifacts often prevalent in dense convolutional operations. The practical implementation of the pixel-sampling Transformer involves selecting $m$ random pixels, denoted $\{x_{i}\}_{i=1}^{m}$, from the valid (masked) region of the input image. For each sampled pixel $x_{i}$, its associated features $F_{\text{enc}}^{f}(x_{i})\in\mathbb{R}^{C}$ are obtained. These features $F_{\text{enc}}^{f}(x_{i})$ are then combined through element-wise addition with $I^{\prime}_{f}(x_{i})\in\mathbb{R}^{C}$. The term $I^{\prime}_{f}(x_{i})$ represents a high-dimensional projection of the raw pixel observations $I_{f}(x_{i})\in\mathbb{R}^{3}$, and this projection is performed by a two-layer MLP with the objective of enhancing the representational power of these raw observations by mapping them to this higher-dimensional space. Notably, our strategy of first projecting the raw observations to $\mathbb{R}^{C}$ and then performing addition differs from SDM UniPS Ikehata (2023), which typically employs direct concatenation of features and the raw observations. These added per-pixel features $F_{\text{add}}^{f}(x_{i})$ are subsequently condensed into compact descriptors $A(x_{i})$ by employing Pooling by Multi-head Attention (PMA) Lee et al. (2019). The resulting collection of $m$ descriptors, $A(x_{i})_{i=1}^{m}$, is then fed into a Transformer network. The processing for each of the $m$ sampled points involves applying frame attention and light-axis attention to aggregate non-local context and cross-image information. Following this, a two-layer MLP is utilized to predict the surface normal vector for each location. These sparsely predicted normals are then systematically merged—for instance, through spatial interpolation or a dedicated upsampling module—to reconstruct the full-resolution surface normal map corresponding to the original input image dimensions. In essence, the pixel-sampling Transformer facilitates the modeling of robust non-local dependencies with notable computational efficiency, while concurrently preserving fine details in the output normal map. This makes the approach particularly well-suited for physics-based vision tasks that involve high-resolution imagery.
@@
In this part, we elaborate on the composition of our total training loss and the design of the respective weights for its constituent components. Based on the definitions provided in Eq. 4, Eq. 1 and Eq. A1, the overall training loss $\mathcal{L}$ for our LINO UniPS method can be expressed as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}+\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$
Let us define the confidence-weighted normal reconstruction loss as $\mathcal{L}_{\text{conf}}$ and the normal gradient supervision loss as $\mathcal{L}_{\text{g}}$. The overall training loss $\mathcal{L}$ can then be expressed as:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}+\lambda_{4}\mathcal{L}_{\text{conf}}+\lambda_{5}\mathcal{L}_{\text{g}},$
@@
A primary motivation for our LINO UniPS framework is to advance beyond prior Universal PS methods by specifically improving the handling of challenging high-frequency regions. We identify these regions based on large magnitudes of the surface normal gradients, as these directly reflect geometric complexity. We deliberately avoid using gradients derived from the input RGB multi-light images as the primary criterion for this identification. The rationale is that while RGB gradients are indeed large in areas of intricate geometric detail, they can also exhibit high magnitudes in regions with significant basecolor variations, which do not necessarily correspond to the geometric high-frequency features we aim to emphasize and reconstruct accurately. Consequently, our methodology incorporates a loss function that is directly informed by surface normal gradients, rather than relying on a naive MSE.
@@
$\lambda_{1}=\frac{0.1}{(\mathcal{L}_{\text{env}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{2}=\frac{0.1}{(\mathcal{L}_{\text{point}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{3}=\frac{0.1}{(\mathcal{L}_{\text{direction}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{4}=1,\lambda_{5}=\frac{0.1}{(\mathcal{L}_{\text{g}}/\mathcal{L}_{\text{conf}})_{\text{sg}}}$
@@
Our decision to set $\lambda_{4}$ to $1$ is because $\mathcal{L}_{\text{conf}}$ serves as the principal component in our overall loss function. The remaining auxiliary losses are then scaled using the adaptive weighting mechanism detailed in Eq. A1.1.7. This mechanism constrains their magnitudes to $0.1$ times that of the primary loss’s detached value, $(\mathcal{L}_{\text{conf}})_{\text{sg}}$, while still allowing their gradients to backpropagate fully. Such a strategy effectively positions these auxiliary losses to act as regularizers to the main learning objective, rather than allowing disparate loss magnitudes to vie for dominance and potentially destabilize training. This controlled weighting is crucial for ensuring stable and efficient training of LINO UniPS, mitigating issues such as excessively slow convergence or even training failure that can arise from an unbalanced multi-term loss function.
A1.2 PBR Materials Prediction
The intrinsic properties of a surface (albedo, metallic, normal, and roughness) are fundamentally entangled, as they are all governed by the underlying surface geometry and material composition. Consequently, our LINO UniPS architecture, designed for normal recovery, can be naturally extended to jointly estimate a full set of PBR material parameters. This extension requires only finetuning on an appropriate dataset with a modified loss function.
A1.2.1 Implementation Details
Our PS-Verse dataset was created with this purpose in mind; alongside surface normals, we also rendered ground truth maps for albedo (3-channel RGB), metallic (1-channel), and roughness (1-channel). A sample of this data is shown in Sec. A1.3.
Starting with our LINO UniPS model pre-trained for normal estimation, we finetune it on this complete PBR dataset for approximately 10 epochs until convergence. This process enables the model to simultaneously predict all four properties from the same input images. For the loss function, we augment the original loss $\mathcal{L}$ with a weighted MSE term for the material maps. The new loss is defined as:
$\mathcal{L_{\text{PBR}}}=\mathcal{L}+\lambda_{6}\mathcal{L}_{\text{a}}+\lambda_{7}\mathcal{L}_{\text{m}}+\lambda_{8}\mathcal{L}_{\text{r}}$
where $\mathcal{L}_{\text{a}},\mathcal{L}_{\text{m}},\mathcal{L}_{\text{r}}$ are the MSE between ground truth and predicted albedo, metallic, and roughness maps, respectively. The coefficients $\lambda_{6},\lambda_{7},\lambda_{8}$ are their corresponding loss weights, which are defined as:
$\lambda_{1}=\frac{1}{(\mathcal{L}_{\text{a}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{2}=\frac{1}{(\mathcal{L}_{\text{m}}/\mathcal{L}_{\text{conf}})_{\text{sg}}},\lambda_{3}=\frac{1}{(\mathcal{L}_{\text{r}}/\mathcal{L}_{\text{conf}})_{\text{sg}}}$
A1.2.2 Experimental Setup
Baselines: Besides the Universal PS methods mentioned in the main paper (UniPS Ikehata (2022), SDM UniPS Ikehata (2023), Uni MS-PS Hardy et al. (2024)), we added one more baseline, IDArb Li et al. (2025). It should be noted that we chose the BRDF version of SDM UniPS, which can predict PBR materials. And IDArb can also predict PBR materials. In contrast, UniPS and Uni MS-PS can only predict normals.
Evaluation Metrics: We evaluate our results using three metrics. For surface normal accuracy, we measure the Mean Angular Error (MAE, in degrees, $\downarrow$), consistent with our main experiments. For the predicted albedo, metallic, and roughness maps, we assess their quality via the Peak Signal-to-Noise Ratio (PSNR, in dB, $\uparrow$) and the Structural Similarity Index (SSIM $\uparrow$).
Evaluation Dataset: The evaluation is conducted on our PS-Verse Testdata, which contains ground truth for all four intrinsic properties.
A1.2.3 Experiment Results
The quantitative results for PBR material estimation are presented in Tab. A1. The results clearly show that our PBR-version LINO UniPS achieves SOTA performance across all predicted maps, including normal, albedo, metallic, and roughness. This comprehensive superiority further validates the effectiveness of our proposed architecture. Furthermore, we observe interesting trade-offs among the baseline methods. While IDArb underperforms all other baselines on the normal prediction task, it surpasses SDM UniPS in estimating the other PBR properties (albedo, metallic, and roughness).
Fig. A1 presents the qualitative results on our PS-Verse Testdata, where our LINO UniPS predictions demonstrate the highest visual fidelity to the ground truth for all properties. In contrast, the baselines exhibit noticeable artifacts. For complex normal estimation, most methods struggle, and IDArb fails completely by predicting a flat surface. For the other material maps, SDM UniPS tends to bake in RGB texture details (e.g., the barrel pattern), while IDArb fails to disentangle complex shading, particularly in shadowed regions like on the soundbox. These visual comparisons further highlight the robustness and superior disentanglement capabilities of our approach.
A1.3 Dataset Analysis and Presentation